OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, 1) and (–2, 5)? A. 4/5 B. 5/4 C. -5/4 D. -4/5

OpenStudy (raden):

formula : slope (m) = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Ik how to do it! I just never get the right answer!!

OpenStudy (anonymous):

Well then i guess i dont how to do t

OpenStudy (raden):

(x1,y1) = (3,1) -----> x1=3 and y1=1 (x2,y2) = (-2,5) -----> x2 = -2 and y2 = 5 u just put of them to formula above

OpenStudy (raden):

m = (y2-y1)/(x2-x1) = (5-1)/(-2-3) = ...

OpenStudy (anonymous):

i got -5

OpenStudy (anonymous):

hello?????

OpenStudy (raden):

5-1 = 4 -2-3 = -5 so, it is 4/-5 or -4/5

OpenStudy (raden):

no options if u got -5 :)